A sample of 300 babies born at a hospital is surveyed. 85% are born on or after their due date. What is the margin of error at a 90% confidence error? 

  • ±2.1% 

  • ±4.0% 

  • ±3.4% 

  • ±5.3%
Answer #1

1. z value for 90% CI is 1.645 as P(-1.645

n=300 and p=0.85

So Margin of Error is E=z*\sqrt{\frac{pq}{n}}=1.65*\sqrt{\frac{0.85*0.15}{300}}=0.034=3.4\%

Hence correct answer here is \pm3.4\%